# A Survey of Large Language Models in Medicine: Progress, Application, and Challenge

### v4

@@ 1 Introduction
Although existing medical LLMs have achieved promising results, there are some key issues in their development and application that need to be addressed. First, many of these models primarily focus on medical dialogue and medical question answering tasks, but their practical utility in clinical practice is often overlooked thirunavukarasu2023large . Recent research and reviews thirunavukarasu2023large ; omiye2024large ; clusmann2023future have begun to explore the potential of medical LLMs in different clinical scenarios, including Electronic Health Records (EHRs) yang2022large , discharge summary generation patel2023chatgpt , health education abd2023large , and care planning medpalm2 . However, they primarily focus on presenting clinical applications of LLMs, especially online commercial LLMs like ChatGPT (including GPT-3.5 and GPT-4 gpt-4 ), without providing practical guidelines for the development of medical LLMs. Besides, they mainly perform case studies to conduct the human evaluation on a small number of samples, thus lacking evaluation datasets for assessing model performance in clinical scenarios. Second, most existing medical LLMs report their performances mainly on answering medical questions, neglecting other biomedical domains, such as medical language understanding and generation. These research gaps motivate this review that offers a comprehensive review of the development of LLMs and their applications in medicine. We aim to cover topics on existing medical LLMs, various medical tasks, clinical applications, and arising challenges.
@@
For the first question, we summarize the principles of existing medical LLMs, detailing their basic structures, the number of parameters, and the data used for model development. Additionally, we provide insight into the construction process of these models. This information is valuable for researchers and medical practitioners looking to build their own medical LLMs tailored to specific needs, such as computational limits, private data, and local knowledge bases. For the second question, we have conducted an extensive review on the performances of existing medical LLMs across ten biomedical NLP (discriminative and generative) tasks. This analysis allows us to understand how medical LLMs outperform traditional medical AI models in different aspects. By showcasing their abilities, we aim to clarify the strengths that medical LLMs bring to the table when deployed in clinical settings. The third question focuses on the practical application of medical LLMs in clinical settings. We provide guidelines for seven clinical application scenarios, offering specific implementations of medical LLMs and highlighting which abilities are used for each scenario. The fourth question emphasizes the challenges that must be overcome when deploying medical LLMs in clinical practice. These challenges include hallucination (i.e. generation of coherent and contextually relevant but factually incorrect outputs) umapathi2023med ; rawte2023survey , ethical, legal, and safety concerns chatgptconcernmed . We also advocate a broader evaluation of medical LLMs, such as trustworthiness xie2023faithful for ensuring their responsible and effective use in clinical settings. For the final question, we look into future directions for developing medical LLMs, providing a guide for researchers and practitioners looking to advance this field and maximize the potential of medical LLMs in medicine.
{forest} for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=edge midpoint, , [Medical Large Language Models, fill=gray!45, parent [ Principles
§ 2, for tree= top_class [Pipeline, for tree=fill=red!45,encoder [Pre-training from Scratch, encoder_more [BioBERTlee2020biobert ; PubMedBERTgu2021domain ; ClinicalBERTalsentzer2019publicly et al., encoder_work] ] [Fine-tuning General LLMs, encoder_more [DoctorGLMxiong2023doctorglm ; HuatuoGPTzhang2023huatuogpt ; LLaVA-Medli2023llavamed et al., encoder_work] ] [Prompting General LLMs, encoder_more [DeID-GPTliu2023deid ; ChatCADwang2023chatcad ; Dr.Knowsgao2023leveraging ; MedPaLMmedpalm et al., encoder_work] ] ] [Data, for tree=fill=green!45,generation [Medical Knowledge Bases, generation_more [DrugBank; NHS Health; NHS Medicine et al., generation_work ] ] [Pre-training Data, generation_more [PubMedPubMed ; MIMIC-IIIjohnson2016mimic ; MedDialogchen2020meddialog et al., generation_work] ] [Fine-tuning Data, generation_more [Qilin-Medye2023qilin ; BianqueCorpuschen2023bianque ; MD-EHRwang2023clinicalgpt et al., generation_work] ] ] ] [, for tree= top_class [Downstream
§ 3, for tree=fill=pink!45,model [Discriminative Tasks, model_wide [ Question Answering; Entity Extraction; Relation Extraction; Text Classification; Natural Language Inference; Semantic Textual Similarity; Information Retrieval et al., model_work ] ] [Generative Tasks, model_wide [Text Generation; Text Summarization; Text Simplification et al., model_work] ] [Performance Comparisons, model_wide [GPT-3.5-turbo; GPT-4; SOTA Task-specific Fine-tuned Models; Human Experts, model_work] ] ] ] [, for tree= top_class [Applications
§ 4, for tree=fill=yellow!45,gpa [Medical Diagnosis, gpa_wide [Chatcadwang2023chatcad ; LLM4medicinethirunavukarasu2023large et al., gpa_work ] ] [Formatting and ICD Coding, gpa_wide [PLM-ICDhuang2022plm ; LLM-ICDong2023applying ., gpa_work ] ] [Clinical Report Generation, gpa_wide [GPT-4V for Medicalwu2023can ; Qilin-Med-VLliu2023qilin et al., gpa_work ] ] [Medical Education, gpa_wide [LLM for medical educationabd2023large et al., gpa_work ] ] [Medical Robotics, gpa_wide [GRIDni2023grid ; Nested -UISxia2023nested et al., gpa_work ] ] [Medical Language Translation, gpa_wide [Advant GLM for Medical Educationkarabacak2023advent et al., gpa_work ] ] [Mental Health Support, gpa_wide [ChatCounselorliu2023chatcounselor et al., gpa_work ] ] ] ] [, for tree= top_class [Challenges
§ 5, for tree=fill=red!45, finetuning [Hallucination, finetuning_wide [Med-haltumapathi2023med ; Selfcheckgptmanakul2023selfcheckgpt et al;, finetuning_work ] ] [Lack of Evaluation Benchmarks and Metrics, finetuning_wide [Truthfulqatruthfulqa ; HaluEvalhalueval et al., finetuning_work ] ] [Domain Data Limitations, finetuning_wide [Model Dementiashumailov2023model et al., finetuning_work ] ] [New Knowledge Adaptation, finetuning_wide [EditingLMMhoelscher2023detecting et al., finetuning_work ] ] [Behavior Alignment, finetuning_wide [Webgptwebgpt ; Chain of Hindsightchainofhindsight et al., finetuning_work ] ] [Ethical & Legal and Safety Concerns, finetuning_wide [Jailbrokenchatgptjailbreak ; Jailpromptshen2023anything et al., finetuning_work ] ] ] ] [, for tree= top_class [Future Directions
§ 6, for tree=fill=yellow!45,inference [New Benchmarks, inference_wide [Chen et al.chen2023extensive ; Chen et al.chen2023large , inference_work ] ] [Multi-modal LLM, inference_wide [Opportunitieshuang2023chatgpt , MedPaLM Mtu2023towards , LLaVA-Medli2023llavamed , Visual Med-AlpacaVisual-Med-Alpaca , Med-Flamingomoor2023medflamingo , Qilin-Med-VLliu2023qilin , inference_work ] ] [Medical Agents, inference_wide [Medagentstang2023medagents , inference_work ] ] [Interdisciplinary Collaborations, inference_wide [Collaborationthirunavukarasu2023large ; omiye2024large ; clusmann2023future ; Malpractice Riskmello2023chatgpt et al., inference_work ] ] ] ] ]
By addressing these questions and providing a holistic perspective on medical LLMs, we hope to promote further research and exploration in this interdisciplinary field. With the rapid development of foundation models, the LLMs could significantly improve future clinical practice and medical discoveries for the benefit of society. It requires sustained interdisciplinary collaboration between clinicians and AI researchers, doctor-in-the-loop, and human-centered design, to address existing challenges with applying LLMs in the medical domain. The co-development of the collaboration for appropriate training data, benchmarks, metrics, and deployment strategies could enable faster and more responsible implementation of medical large language models than the current practice.
2 The Principles of Medical Large Language Models
Box 1 and Table LABEL:table:llm_summary briefly introduce the background of general LLMs zhao2023survey , e.g., GPT-4 gpt-4 . Table 1 summarizes the currently available medical LLMs according to their model development. Existing medical LLMs are mainly pre-trained from scratch, fine-tuned from existing general LLMs, or directly obtained through prompting to align the general LLMs to the medical domain. Therefore, we introduce the principles of medical LLMs in terms of these three methods: pre-training, fine-tuning, and prompting. Meanwhile, we further summarize the medical LLMs according to their model architectures in Figure 2.
2.1 Pre-training
To meet the needs of the medical domain, pre-training medical LLMs typically involve refining the following commonly used training objectives in general LLMs: masked language modeling, next sentence prediction, and next token prediction (Please see Box 1 for an introduction of these three pre-training objectives). For example. BERT-series models (e.g., BioBERT lee2020biobert , PubMedBERT gu2021domain , ClinicalBERT alsentzer2019publicly , and GatorTron yang2022large ) mainly adopt the masked language modeling and the next sentence prediction for pre-training; GPT-series models (e.g., BioGPT luo2022biogpt , and GatorTronGPT peng2023study ) mainly adopt the next token prediction for pre-training. After pre-training, medical LLMs can learn rich medical knowledge that can be leveraged to achieve strong performance on different medical tasks
@@ 2.1 Pre-training
Table 1: Summary of existing medical LLMs, in terms of their model development, the number (#) of parameters, the scale of pre-training/fine-tuning data, and the data source. M: million, B: billion.
@@
Medical-domain LLMs (Sec. 2) Pre-training (Sec. 2.1) BioBERT lee2020biobert ; raj2019saama 110M 18B tokens PubMed PubMed
PubMedBERT gu2021domain 110M/340M 3.2B tokens PubMed PubMed
@@
BioMedLM venigalla2022biomedlm 2.7B 110GB PubMed pile
@@
GatorTron yang2022gatortron ; yang2022large 8.9B
>82B tokens
6B tokens
@@
EHRs yang2022large
PubMed PubMed
@@
MEDITRON chen2023meditron 70B 48.1B tokens
PubMed PubMed
Clinical Guidelines chen2023meditron
@@
VariousMedQA zhang2018multi ; medqa
@@
HealthCareMagic healthcaremagic
iCliniq icliniq
@@
Books+Literature lo2019s2orc
MedC-I wu2023pmc
@@
eICU-CRD pollard2018eicu
MIMIC-IV johnson2020mimiciv
@@
ShareGPT sharegpt
PubMed PubMed
@@
BioMedGPT luo2023biomedgpt 7B >26B tokens S2ORC lo2019s2orc
@@
Dr. Knows gao2023leveraging ChatGPT Zero-shot Prompting UMLS bodenreider2004unified ; lindberg1993unified
@@ 2.2 Fine-tuning
It is high-cost and time-consuming to train a medical LLM from scratch, due to its requirement of substantial (e.g. several days or even weeks) computational power and manual labor. One solution is to fine-tune the general LLMs with medical data, and researchers have proposed different fine-tuning methods medpalm2 ; han2023medalpaca ; toma2023clinical for learning domain-specific medical knowledge and obtain medical LLMs. Current fine-tuning methods include Supervised Fine-Tuning (SFT), Instruction Fine-Tuning (IFT), and Parameter-Efficient Tuning. The resulting fine-tuned medical LLMs are summarized in Table 1.
@@
Parameter-Efficient Tuning
aims to substantially reduce computational and memory requirements for fine-tuning general LLMs. The main idea is to keep most of the parameters in pre-trained LLMs unchanged, by fine-tuning only the smallest subset of parameters (or additional parameters) in these LLMs. Commonly used parameter-efficient tuning techniques include Low-Rank Adaptation (LoRA) lora , Prefix Tuning li2021prefixtuning , and Adapter Tuning liu2022p ; liu2021p .
In contrast to fine-tuning full-rank weight matrices, 1) LoRA preserves the parameters of the original LLMs and only adds trainable low-rank matrices into the self-attention module of each Transformer layer lora . Therefore, LoRA can significantly reduce the number of trainable parameters and improve the efficiency of fine-tuning, while still enabling the fine-tuned LLM to capture effectively the characteristics of the downstream tasks. 2) Prefix Tuning takes a different approach from LoRA by adding a small set of continuous task-specific vectors (i.e. “prefixes”) to the input of each Transformer layer li2021prefixtuning ; zhao2023survey . These prefixes serve as the additional context to guide the generation of the model without changing the original pre-trained parameter weights. 3) Adapter Tuning involves introducing small neural network modules, known as adapters, into each Transformer layer of the pre-trained LLMs houlsby2019parameter . These adapters are fine-tuned while keeping the original model parameters frozen houlsby2019parameter , thus allowing for flexible and efficient fine-tuning. The number of trainable parameters introduced by adapters is relatively small, yet they enable the LLMs to adapt to downstream scenarios or tasks effectively.
In general, parameter-efficient tuning is valuable for developing LLMs that meet unique needs in specific (e.g. medical) domains, due to its ability to reduce computational demands while maintaining the model performance. For example, medical LLMs DoctorGLM xiong2023doctorglm , MedAlpaca han2023medalpaca , Baize-Healthcare xu2023baize , Zhongjing yang2023zhongjing , CPLLM shoham2023cpllm , and Clinical Camel toma2023clinical adopted the LoRA lora to perform parameter-efficient fine-tuning to efficiently align the general LLMs to the medical domain.

### v5

@@ 1 Introduction
Although existing medical LLMs have achieved promising results, there are some key issues in their development and application that need to be addressed. First, many of these models primarily focus on medical dialogue and medical question-answering tasks, but their practical utility in clinical practice is often overlooked thirunavukarasu2023large. Recent research and reviews thirunavukarasu2023large; omiye2024large; clusmann2023future have begun to explore the potential of medical LLMs in different clinical scenarios, including Electronic Health Records (EHRs) yang2022large, discharge summary generation patel2023chatgpt, health education abd2023large, and care planning medpalm2. However, they primarily focus on presenting clinical applications of LLMs, especially online commercial LLMs like ChatGPT (including GPT-3.5 and GPT-4 gpt-4), without providing practical guidelines for the development of medical LLMs. Besides, they mainly perform case studies to conduct the human evaluation on a small number of samples, thus lacking evaluation datasets for assessing model performance in clinical scenarios. Second, most existing medical LLMs report their performances mainly on answering medical questions, neglecting other biomedical domains, such as medical language understanding and generation. These research gaps motivate this review which offers a comprehensive review of the development of LLMs and their applications in medicine. We aim to cover topics on existing medical LLMs, various medical tasks, clinical applications, and arising challenges.
@@
For the first question, we analyze the foundational principles underpinning current medical LLMs, providing detailed descriptions of their architecture, parameter scales, and the datasets used during their development. This exposition aims to serve as a valuable resource for researchers and clinicians designing medical LLMs tailored to specific requirements, such as computational constraints, data privacy concerns, and the integration of local knowledge bases. For the second question, we evaluate the performance of medical LLMs across ten biomedical NLP tasks, encompassing both discriminative and generative tasks. This comparative analysis elucidates how these models outperform traditional AI approaches, offering insights into the specific capabilities that render LLMs effective in clinical environments. The third question, the practical deployment of medical LLMs in clinical settings, is explored through the development of guidelines tailored for seven distinct clinical application scenarios. This section outlines practical implementations, emphasizing specific functionalities of medical LLMs that are leveraged in each scenario. The fourth question emphasizes addressing the challenges associated with the clinical deployment of medical LLMs, such as the risk of generating factually inaccurate yet plausible outputs (hallucination), and the ethical, legal, and safety implications. Citing recent studies, we argue for a comprehensive evaluation framework that assesses the trustworthiness of medical LLMs to ensure their responsible and effective utilization in healthcare. For the last question, we propose future research directions to advance the medical LLMs field. This includes fostering interdisciplinary collaboration between AI specialists and medical professionals, advocating for a ’doctor-in-the-loop’ approach, and emphasizing human-centered design principles.
By establishing robust training data, benchmarks, metrics, and deployment strategies through co-development efforts, we aim to accelerate responsible and efficacious integration of medical LLMs into clinical practice. This study therefore seeks to stimulate continued research and development in this interdisciplinary field, with the objective of realizing the profound potential of medical LLMs in enhancing clinical practice and advancing medical science for for the betterment of society.
BOX 1: Background of Large Language Models (LLMs)
  The impressive performance of LLMs can be attributed to Transformer-based language models, large-scale pre-training, and scaling laws. Language Models    A language model bengio2000neural; mikolov2010recurrent; sundermeyer2015feedforward is a probabilistic model that models the joint probability distribution of tokens (meaningful units of text, such as words or subwords or morphemes) in a sequence, i.e., the probabilities of how words and phrases are used in sequences. Therefore, it can predict the likelihood of a sequence of tokens given the previous tokens, which can be used to predict the next token in a sequence or to generate new sequences. The Transformer architecture    The recurrent neural network (RNN) Hochreiter1997LSTM; mikolov2010recurrent has been widely used for language modeling by processing tokens sequentially and maintaining a vector named hidden state that encodes the context of previous tokens. Nonetheless, sequential processing makes it unsuitable for parallel training and limits its ability to capture long-range dependencies, making it computationally expensive and hindering its learning ability for long sequences. The strength of the Transformer transformer lies in its fully attentive mechanism, which relies exclusively on the attention mechanism and eliminates the need for recurrence. When processing each token, the attention mechanism computes a weighted sum of the other input tokens, where the weights are determined by the relevance between each input token and the current token. It allows the model to adaptively focus on different parts of the sequence to effectively learn the joint probability distribution of tokens. Therefore, Transformer not only enables efficient modeling of long-text but also allows highly paralleled training devlin2018bert, thus reducing training costs. They make the Transformer highly scalable, and therefore it is efficient to obtain LLMs through the large-scale pre-training strategy. Large-scale Pre-training    The LLMs are trained on massive corpora of unlabeled texts (e.g., CommonCrawl, Wiki, and Books) to learn rich linguistic knowledge and language patterns. The common training objectives are masked language modeling (MLM) and next token prediction (NTP). In MLM, a portion of the input text is masked, and the model is tasked with predicting the masked text based on the remaining unmasked context, encouraging the model to capture the semantic and syntactic relationships between tokens devlin2018bert; NTP is another common training objective, where the model is required to predict the next token in a sequence given the previous tokens. It helps the model to predict the next token GPT-3. Scaling Laws    LLMs are the scaled-up versions of Transformer architecture transformer with increased numbers of Transformer layers, model parameters, and volume of pre-training data. The “scaling laws” kaplan2020scaling; hoffmann2022training predict how much improvement can be expected in a model’s performance as its size increases (in terms of parameters, layers, data, or the amount of training computed). The scaling laws proposed by OpenAI kaplan2020scaling show that to achieve optimal model performance, the budget allocation for model size should be larger than the data. The scaling laws proposed by Google DeepMind hoffmann2022training show that both model and data sizes should be increased in equal scales. The scaling laws guide researchers to allocate resources and anticipate the benefits of scaling models. General Large Language Models    Existing general LLMs can be divided into three categories based on their architecture (Table 1). Encoder-only LLMs    consisting of a stack of Transformer encoder layers, employ a bidirectional training strategy that allows them to integrate context from both the left and the right of a given token in the input sequence. This bi-directionality enables the models to achieve a deep understanding of the input sentences devlin2018bert. Therefore, encoder-only LLMs are particularly suitable for language understanding tasks (e.g., sentiment analysis document classification) where the full context of the input is essential for accurate predictions. BERT devlin2018bert and DeBERTa he2021deberta are the representative encoder-only LLMs. Decoder-only LLMs    utilize a stack of Transformer decoder layers and are characterized by their uni-directional (left-to-right) processing of text, enabling them to generate language sequentially. This architecture is trained unidirectionally using the next token prediction training objective to predict the next token in a sequence, given all the previous tokens. After training, the decoder-only LLMs generate sequences autoregressively (i.e. token-by-token). The examples are the GPT-series developed by OpenAI GPT-3; gpt-4, the LLaMA-series developed by Meta llama; touvron2023llama2, and the PaLM paLM and Bard (Gemini) google2023bard developed by Google. Based on the LLaMA model, Alpaca alpaca is fine-tuned with 52k self-instructed data supervision. In addition, Baichuan yang2023baichuan is trained on approximately 1.2 trillion tokens that support bilingual communication in Chinese and English. These LLMs have been used successfully in language generation. Encoder-decoder LLMs    are designed to simultaneously process input sequences and generate output sequences. They consist of a stack of bidirectional Transformer encoder layers followed by a stack of unidirectional Transformer decoder layers. The encoder processes and understands the input sequences, while the decoder generates the output sequences du2022glm; zeng2022glm; chung2022scaling. Representative examples of encoder-decoder LLMs include Flan-T5 joseph2023multilingual, and ChatGLM du2022glm; zeng2022glm. Specifically, ChatGLM du2022glm; zeng2022glm has 6.2B parameters and is a conversational open-source LLM specially optimized for Chinese to support Chinese-English bilingual question-answering.
Table 1: Summary of existing general (large) language models, their underlying structures, numbers of parameters, and datasets used for model training. Column “# params” shows the number of parameters, M: million, B: billion. Domains Model Structures Models # Params Pre-train Data Scale General-domain (Large) Language Models Encoder-only BERT devlin2018bert 110M/340M 3.3B tokens RoBERTa liu2019roberta 355M 161GB DeBERTahe2021deberta 1.5B 160GB Decoder-only GPT-2gpt-2 1.5B 40GB Vicunavicuna 7B/13B LLaMA + 70K dialogues Alpacaalpaca 7B/13B LLaMA+ 52K IFT Mistraljiang2023mistral 7B - LLaMA llama 7B/13B/33B/65B 1.4T tokens LLaMA-2 touvron2023llama2 7B/13B/34B/70B 2T tokens LLaMA-3 meta_llama3 8B/70B 15T tokens GPT-3GPT-3 6.7B/13B/175B 300B tokens Qwenbai2023qwen 1.8B/7B/14B/72B 3T tokens PaLM paLM 8B/62B/540B 780B tokens FLAN-PaLM chung2022scaling 540B - Gemini (Bard) google2023bard - - GPT-3.5 instructgpt - - GPT-4gpt-4 - - Claude-3Claude2024 - - Encoder-Decoder BART lewis2019bart 140M/400M 160GB ChatGLM du2022glm; zeng2022glm 6.2B 1T tokens T5 joseph2023multilingual 11B 1T tokens FLAN-T5 chung2022scaling 3B/11B 780B tokens UL2 tay2022ul2 19.5B 1T tokens GLM zeng2022glm 130B 400B tokens
@@
Table 2: Summary of existing medical-domain LLMs, in terms of their model development, the number of parameters (# params), the scale of pre-training/fine-tuning data, and the data source. M: million, B: billion.
@@
Medical-domain LLMs (Sec. 2) Pre-training (Sec. 2.1) BioBERT lee2020biobert 110M 18B tokens PubMed PubMed+PMC PMC
PubMedBERT gu2021domain 110M/340M 3.2B tokens PubMed PubMed+PMC PMC
@@
BioM-ELECTRA alrowili2021large 110M/335M - PubMed PubMed
BioMed-RoBERTa domains 125M 7.55B tokens S2ORC lo2019s2orc
BioLinkBERT yasunaga2022linkbert 110M/340M 21GB PubMed PubMed
SciFive phan2021scifive 220M/770M - PubMed PubMed+PMC PMC
ClinicalT5 lu2022clinicalt5 220M/770M 2M clinical notes MIMIC-III johnson2016mimic
BlueBERT peng2019transfer; mutinda2020detecting; mahajan2020identification 110M/340M >4.5B tokens
PubMedPubMed+MIMIC-IIIjohnson2016mimic
@@
BioMedLM venigalla2022biomedlm 2.7B 110GB Pile pile
@@
GatorTron yang2022large 8.9B
>82B tokens+6B tokens
@@
EHRsyang2022large+PubMedPubMed
@@
VariousMedQA medqa
@@
HealthCareMagic healthcaremagic+iCliniq icliniq
@@
Books+Literature lo2019s2orc+MedC-I wu2023pmc
@@
eICU-CRD pollard2018eicu+MIMIC-IV johnson2020mimiciv
OpenBioLLM OpenBioLLMs 8B/70B -
MEDITRON chen2023meditron; bosselut2024meditron 7B/70B 48.1B tokens
PubMed PubMed+Guidelines chen2023meditron
@@
ShareGPTsharegpt+PubMedPubMed
@@
Med-Gemini yang2024advancing; saab2024capabilities - -
MedQA-R&RS saab2024capabilities++MultiMedQA medpalm2
+MIMIC-III johnson2016mimic+MultiMedBench tanno2024consensus
@@
Dr. Knows gao2023leveraging ChatGPT Zero-shot Prompting UMLS bodenreider2004unified
@@
Chat-Orthopedist shi2023retrieval ChatGPT
Retrieval-Augmented
Generation (RAG)
PubMed+GuidelinesSRS +
UpToDateUpToDate+DynameDynamed
QA-RAG kim2024rag ChatGPT RAG FDA QA kim2024rag
Almanac zakka2024almanac ChatGPT RAG & CoT Clinical QA zakka2024almanac
@@
2 The Principles of Medical Large Language Models
Box 1 and Table 1 briefly introduce the background of general LLMs zhao2023survey, e.g., GPT-4 gpt-4. Table 2 summarizes the currently available medical LLMs according to their model development. Existing medical LLMs are mainly pre-trained from scratch, fine-tuned from existing general LLMs, or directly obtained through prompting to align the general LLMs to the medical domain. Therefore, we introduce the principles of medical LLMs in terms of these three methods: pre-training, fine-tuning, and prompting. Meanwhile, we further summarize the medical LLMs according to their model architectures in Figure 2.
2.1 Pre-training
To meet the needs of the medical domain, pre-training medical LLMs typically involve refining the following commonly used training objectives in general LLMs: masked language modeling, next sentence prediction, and next token prediction (Please see Box 1 for an introduction of these three pre-training objectives). For example. BERT-series models (e.g., BioBERT lee2020biobert, PubMedBERT gu2021domain, ClinicalBERT alsentzer2019publicly, and GatorTron yang2022large) mainly adopt the masked language modeling and the next sentence prediction for pre-training; GPT-series models (e.g., BioGPT luo2022biogpt, and GatorTronGPT peng2023study) mainly adopt the next token prediction for pre-training. It is worth mentioning that BERT-like Medical LLMs (e.g. BioBERT lee2020biobert, PubMedBERT gu2021domain, Clinical BERT alsentzer2019publicly) are originally derived from the general domain BERT or RoBERTa models. To clarify the differences between different models, in our Table 2, we only show the data source used to further construct medical LLMs. After pre-training, medical LLMs can learn rich medical knowledge that can be leveraged to achieve strong performance on different medical tasks.
@@ 2.2 Fine-tuning
It is high-cost and time-consuming to train a medical LLM from scratch, due to its requirement of substantial (e.g. several days or even weeks) computational power and manual labor. One solution is to fine-tune the general LLMs with medical data, and researchers have proposed different fine-tuning methods medpalm2; han2023medalpaca; toma2023clinical for learning domain-specific medical knowledge and obtaining medical LLMs. Current fine-tuning methods include Supervised Fine-Tuning (SFT), Instruction Fine-Tuning (IFT), and Parameter-Efficient Fine-Tuning (PEFT). The resulting fine-tuned medical LLMs are summarized in Table 2.
@@
Parameter-Efficient Fine-Tuning (PEFT)
aims to substantially reduce computational and memory requirements for fine-tuning general LLMs. The main idea is to keep most of the parameters in pre-trained LLMs unchanged, by fine-tuning only the smallest subset of parameters (or additional parameters) in these LLMs. Commonly used PEFT techniques include Low-Rank Adaptation (LoRA) lora, Prefix Tuning li2021prefixtuning, and Adapter Tuning liu2022p; liu2021p.
In contrast to fine-tuning full-rank weight matrices, 1) LoRA preserves the parameters of the original LLMs and only adds trainable low-rank matrices into the self-attention module of each Transformer layer lora. Therefore, LoRA can substantially reduce the number of trainable parameters and improve the efficiency of fine-tuning, while still enabling the fine-tuned LLM to capture effectively the characteristics of the tasks. 2) Prefix Tuning takes a different approach from LoRA by adding a small set of continuous task-specific vectors (i.e. “prefixes”) to the input of each Transformer layer li2021prefixtuning; zhao2023survey. These prefixes serve as the additional context to guide the generation of the model without changing the original pre-trained parameter weights. 3) Adapter Tuning involves introducing small neural network modules, known as adapters, into each Transformer layer of the pre-trained LLMs houlsby2019parameter. These adapters are fine-tuned while keeping the original model parameters frozen houlsby2019parameter, thus allowing for flexible and efficient fine-tuning. The number of trainable parameters introduced by adapters is relatively small, yet they enable the LLMs to adapt to clinical scenarios or tasks effectively.
In general, PEFT is valuable for developing LLMs that meet unique needs in specific (e.g. medical) domains, due to its ability to reduce computational demands while maintaining the model performance. For example, medical LLMs DoctorGLM xiong2023doctorglm, MedAlpaca han2023medalpaca, Baize-Healthcare xu2023baize, Zhongjing yang2023zhongjing, CPLLM shoham2023cpllm, and Clinical Camel toma2023clinical adopted the LoRA lora to perform parameter-efficient fine-tuning to efficiently align the general LLMs to the medical domain.